SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company’s products is grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week, the company prepared 3,500 of these meals using 1,710 direct labor-hours. The company paid its direct labor workers a total of $23,940 for this work, or $14.00 per hour.

According to the standard cost card for this meal, it should require 0.50 direct labor-hours at a cost of $13.50 per hour.

Required:

1. What is the standard labor-hours allowed (SH) to prepare 3,500 meals?

2. What is the standard labor cost allowed (SH × SR) to prepare 3,500 meals?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?
Answer #1

2 No. of meals prepared = 3500 Standard direct labour-hour por meal = 0:50 Standard labour-hours allowed to prepare 3500 mealLaboy rate variance = (sR-AR) * AH = ($13.50 - $14) * 1710 hrs = $ 855 U Labor efficiency variance = (SH-AH) * SR = (1750 - 1